10×121^0.5/100^0.5

​=110/10=11


Please how come you got 12.1 as answer


1
Expert's answer
2022-04-15T16:38:59-0400

"\frac{10\cdot 121^{0.5}}{100^{0.5}}=\frac{10\cdot 11}{10}=11."